# fmeffects: An R Package for Forward Marginal Effects

### v1

@@ Introduction
Let us take a moment to compare the FME plot with the individual conditional expectation (ICE) (Goldstein et al., 2015) and partial dependence (PD) (Friedman, 2001) plot generated by the R package iml (Molnar et al., 2018). The ICE and PD are one of the most popular and established model-agnostic ways to interpret predictive models (Molnar, 2022). The ICE is a local model explanation and represents the prediction for an observation where only the features of interest are varied (in this case only ‘temp’). The PD is the average of ICEs (in the univariate case, the vertical average) and indicates the global, average prediction when a subset of features is varied for all observations. From Fig. 4, we gather that in contrast to FMEs, the ICE and PD plot is more difficult to interpret. Although we can see a rough trajectory of the feature influence on local and average predictions, it is difficult to pinpoint the exact effects of changing ‘temp’ on the prediction for single observations. Furthermore, ICE curves are more likely to be subject to model extrapolations, a result from predicting in areas where the model was not trained on a sufficient amount of data.
@@
Figure 4: An ICE and PD plot for feature ‘temp’ generated by the R package iml. Each black curve (an ICE) represents predictions for a single instance while only feature ‘temp’ varies. The yellow curve (the PD) is the vertical average of ICEs and represents the average, isolated influence of ‘temp’ on predictions. As opposed to FMEs, this plot is substantially more difficult to interpret.
In many applications, we are interested in interactions of features on the prediction. We diagnosed a univariate effect of ‘temp’ on the predicted amount of bike rentals, but effects may be enhanced or mitigated when other features are varied at the same time. For instance, humidity has a considerable impact on the effects of heat (Russo et al., 2017). We evaluate an increase in temperature by 5 degrees Celsius and a simultaneous drop in humidity by 10 percentage points (see Fig. 5). For a bivariate change in feature values, the two arrows depict the direction and magnitude of the feature change in the respective variable. Each point is located at the observation’s observed feature values in the provided data set. Its color indicates the FME associated with the bivariate feature change. An increase in the outside temperature by 5 degrees Celsius and a drop in humidity by 10 percentage points is associated with an AME of 19.14. The isolated univariate effect of increasing the outside temperature is boosted by the simultaneous drop in humidity.
@@
> effects.bivariate = fme(model = forest,+               data = bikes,+              target = "count",+              feature = c("temp", "humidity"),+              step.size = c(5, -0.1),+              ep.method = "envelope")> summary(effects.bivariate)
@@
Forward Marginal Effects ObjectStep type:  numericalFeatures & step lengths:  temp, 5  humidity, -0.1Extrapolation point detection:  envelope, EPs: 82 of 727 obs. (11 %)Average Marginal Effect (AME):  19.1418
@@
Figure 5: Visualizing bivariate FMEs for an increase in ‘temp’ by 5 degrees Celsius and a simultaneous drop in ‘humidity’ by 10 percentage points. FMEs are highly heterogeneous. We can see mostly positive effects, especially for observations with combinations of lower ‘temp’ and ‘humidity’ values. For various really warm days, the model even predicts a negative effect.
@@
Let us repeat the same procedure as for univariate feature changes and compare the FME plot to an alternative option, the bivariate PD plot (see Fig. 6). As opposed to the novel visualization with FMEs, the PD plot only visualizes the average, global effect of changing both features on the predicted amount of bike rentals, and it does not inform about the distribution of observed feature values, thus not allowing us to evaluate effects of increasing one feature and decreasing another at the same time. Furthermore, the PD plot - like its univariate counterpart - is more likely to be subject to model extrapolations due to predicting for many feature value combinations the model has not seen during training.
@@
Figure 6: A bivariate PD plot (created via the R package iml), one of the few available visualization tools for interactions. We can see the global interaction between ‘temp’ and ‘humidity’ on the predicted amount of bike rentals. Plugging in larger values for both features ceteris paribus results in more predicted bike rentals on average. As opposed to bivariate FMEs, we cannot investigate multiple local effects, nor can we see the actual distribution of observed feature values, and as a result, we cannot evaluate the effects of increasing one feature and decreasing another at the same time.
@@
Let us now proceed to investigations of non-linearity. Although some sense of non-linearity can be derived from ICE curves (see Fig. 4), it is hard to quantify, and summarizing non-linearity for an ICE with a large variation in the feature of interest would be somewhat meaningless. Furthermore, for bivariate or higher-dimensional changes in feature values, we lose any option for visual diagnoses of non-linearity. In contrast, the NLM can be computed for FMEs with continuous step sizes, regardless of their dimensionality:
@@
> effects.bivariate.nlm = fme(model = forest,+                       data = bikes,+                       target = "count",+                       feature = c("temp", "humidity"),+                       step.size = c(5, -0.1),+                       ep.method = "envelope",+                       compute.nlm = TRUE)> plot(effects.bivariate.nlm, with.nlm = TRUE)
Figure 7: Adding NLM computations to the FME plot. Each dot in the left and right plot represents a single observation and its respective FME and NLM value. We can see both linear (bright, yellow colored NLM) and non-linear FMEs (dark, blue colored NLM).
Fig. 7 simply contrasts each observation’s NLM value with its FME value. In this case, we can see both non-linear FMEs (darker, blue colored NLM values) and linear FMEs (brighter, yellow colored NLM values). This is not surprising, considering that the bike sharing data consists of many features, each being capable to influence the effect of changes in ‘temp’ and ‘humidity’ on the prediction.
@@
So far, we have only evaluated changes in continuous features. In many applications, we are often concerned with switching categories of categorical features, a way of counterfactual thinking inherent to the human thought process. Note that in spite of the allure of switching categories of categorical features, one needs to be aware of potential model extrapolations. To illustrate this, we switch each non-rainy day’s precipitation status to rainfall. Rainfall has an average isolated effect of lowering daily rentals by 57 bikes (see Fig. 8).
@@
> effects.categ = fme(model = forest,+                data = bikes,+                target = "count",+                feature = "weather",+                step.size = "rain")> summary(effects.categ)Forward Marginal Effects ObjectStep type:  categoricalFeature & reference category:  weather, rainExtrapolation point detection:  none, EPs: 0 of 657 obs. (0 %)Average Marginal Effect (AME):  -57.3322> plot(effects.categ)
@@
Figure 8: Distribution of categorical FMEs resulting from switching each non-rainy day’s precipitation status to rain. On average, rainfall lowers predicted bike rentals by 57 bikes per day.
@@
In our examples, we can see highly heterogeneous local effects. The more heterogeneous FMEs are, the less information the AME carries. In many practical applications, we are interested in compactly describing the behavior of the predictive model across the feature space, akin to a beta coefficient in a linear model. This is where regional explanations come into play. We aim to find subgroups with more homogeneous FME values, thereby describing the behavior of the model not in terms of a global average, but in terms of multiple regional averages (cAMEs).
In fmeffects, this can be achieved by further processing the ForwardMarginalEffects object containing FMEs (and optionally NLM values) using the came() function. This returns a Partitioning object (in this case, an object of the class "PartitioningCTREE", a subclass of the abstract class "Partitioning", see later section on Design and options for extensions).
For the univariate change in temperature by 5 degrees Celsius, we decide to search for exactly 3 subgroups [^4] This value is to be set by the user depending on how many regional explanations are to be found. How many subgroups can be found depends on the data and predictive model. (for a description of this algorithm, see the following section on Design and options for extensions). A summary of the created object informs us about the number of observations, cAME, and standard deviation (SD) of FMEs inside the root node and leaf nodes (the found subgroups). We succeed in finding subgroups with lower SDs while maintaining an appropriate sample size. The root node SD of 22.5 can be successfully split down to 21.5, 8.5, and 17.7 inside the subgroups. By visualizing the tree, we can see how the data was partitioned. Lower temperatures are associated with larger increases in predicted bike rentals when the outside temperature is raised. For outside temperatures lower than 17 degrees Celsius and working days, we can observe the largest cAME of 32 additional bike rentals per day and a cAME of 11 additional daily bike rentals for weekends. For outside temperatures above 17 degrees Celsius, there is only a slight effect of 1.7 additional bikes per day:
> subspaces = came(effects = effects, number.partitions = 3)> summary(subspaces)
PartitioningCtree of an FME objectMethod:  partitions = 3   n      cAME   SD(FME) 646 14.484353 22.499156 * 230 32.265791 21.546565 130 11.172507  8.554322 286  1.689979 17.749573---* root node (non-partitioned)AME (Global): 14.4844
@@
Figure 9: Finding subgroups of observations with a decision tree. Each leaf node visualizes one subgroup, the number of observations, the cAME, and the SD of FMEs indicating FME heterogeneity.
@@
When to search for regional explanations thus depends on the heterogeneity of local effects, which can be summarized for the entire model via the ame() function. It uses a default step size of 1 or 0.01 for small feature ranges. For categorical FMEs, it uses every observed category as a reference category. The summary() function prints a compact model summary of each feature, a default step size, the AME, the SD of FMEs, 25% and 75% quantiles of FMEs, as well as the number of observations left after excluding extrapolation points. A large dispersion indicates heterogeneity of FMEs and thus, a small fidelity of the AME and possible benefits from searching for subgroups with varying cAMEs. A different workflow can, therefore, also consist of starting with the table generated by ame() and deciding which feature effects can be described by AMEs, which might be better describable by subgroups and cAMEs, and - if this has been without success - resorting to local model explanations. Recall our example from the previous section on Regional explanations where we split FMEs associated with increasing temperature by 5 degrees Celsius. From the ame() summary, we gather that ‘temp’ has a relatively large SD in relation to its AME, and the interquartile range indicates a large spread of FMEs from -0.47 in the 25% quantile up to 4.74 in the 75% quantile, thus making it a promising candidate to find subgroups with more homogeneous FMEs.
> ame.results = ame(model = forest, data = bikes, target = "count")> summary(ame.results)
@@
Model Summary Using Average Marginal Effects:      feature step size       AME      SD      0.25      0.75   n1      season    spring  -29.3415 30.6189  -38.2318   -6.3963 5482      season    summer    0.5742 21.4782   -8.3499   10.5607 5433      season      fall      9.97 27.4275   -1.9819    30.745 5394      season    winter   13.9861 23.6806    1.0266   23.6969 5515        year         0  -98.6521 67.2816 -156.4247    -19.63 3646        year         1   96.6155 60.1458     22.91  148.1915 3637       month         1    3.9954 13.4096   -1.2981    7.2207 7278     holiday     False   -1.5799 20.5126  -12.5858   14.0711  219     holiday      True  -12.0674 24.2779  -29.8895    6.4375 70610    weekday       Sat  -54.4458 49.1836  -85.8245  -15.5722 62211    weekday       Sun  -78.6803  51.462 -113.7695  -31.2021 62212    weekday       Mon   10.7612 28.3258   -8.7131   30.5779 62313    weekday       Tue   17.9565 24.7443    0.9637   33.5499 62514    weekday       Wed    20.721 22.6949    1.5863   34.3365 62315    weekday       Thu   19.7423 23.9554   -0.2864   34.7076 62416    weekday       Fri    1.6724 35.8582  -23.9194   29.7121 62317 workingday     False -204.2023 90.1427 -256.0286 -141.4579 49618 workingday      True  165.2412 64.8985  123.8688   213.351 23119    weather     clear   25.9084  42.223    3.6162   22.3112 28420    weather     misty    3.3301 33.2101   -9.0864    0.8842 51321    weather      rain  -57.3322 55.5617  -95.7276   -3.8571 65722       temp         1    2.3959  7.2201   -0.4717    4.7369 72723   humidity      0.01   -0.1681  2.4691   -0.2999    0.4329 72724  windspeed         1    0.0198  2.2861   -0.2156    0.2126 727
@@
The fmeffects package is built on a modular design for improved maintainability and future extensions. Fig. 10 provides a visual overview of the core design. Major emphasis is given to the strategy and adapter design patterns (Gamma et al., 1994). Simply put, the strategy pattern decouples source code for algorithm selection at runtime into separate classes. We repeatedly implement this pattern throughout the package by creating abstract classes whose subclasses implement specific functionalities. The adapter design pattern (also referred to as a “wrapper”) creates an interface for communication between two classes.
@@
  • "Predictor": An abstract class that implements the adapter pattern to accommodate future implementations of storing a predictive model. "PredictorMLR3", "PredictorParsnip", and "PredictorCaret" are subclasses that store an mlr3, parsnip (Kuhn and Vaughan, 2023) (part of tidymodels), or caret model object. This allows users of fmeffects to use numerous predictive models such as random forests, gradient boosting, support vector machines, or neural networks. The package can be extended with novel model types by implementing a new subclass that stores the model, data, target, and is able to return predictions.
  • "AverageMarginalEffects": A class to compute AMEs for each feature in the data (or a subset thereof). Internally, a new "ForwardMarginalEffects" object is used to compute FMEs and aggregate them. For convenience, we implement a wrapper function ame() to facilitate object creation and to initiate computations without requiring user input in the form of R6 syntax.
  • "ForwardMarginalEffects": The centerpiece class of the package. It keeps access to a Predictor, stores important information to create FMEs, and after computations are completed, stores results and gives access to visualization methods. For convenience, the wrapper function fme() can be used.
  • "FMEPlot": An abstract class for code decoupling of different plot categories into distinct classes. Subclasses include "FMEPlotUnivariate", "FMEPlotBivariate", "FMEPlotCategorical".
  • "ExtrapolationDetector": Identifies (and excludes) extrapolation points. The current implementation supports the method “envelope”, which excludes points outside of the multivariate envelope of the training data (Scholbeck et al., 2022).
  • "NonLinearityMeasure": For the NLM, we need to approximate three line integrals, for instance via Simpson’s 3/8 rule. The general definition for a univariate function f(x) and integration interval $[a,b]$ corresponds to:
@@
  • "Partitioning": An abstract class, allowing various implementations of finding subgroups for cAMEs. For convenience, the wrapper function came() can be used. The current implementation supports RP via the rpart and partykit (CTREE algorithm) packages (classes "PartitioningRPart" and "PartitioningCTREE"). We believe there are two criteria that should guide this process:
    • FME homogeneity inside each subgroup.
    • The number of subgroups.
    A low number of subgroups is generally preferred. In certain applications, we may want to search for a predefined number of subgroups, akin to the search for a predefined number of clusters in clustering problems. Many RP algorithms do not support the search for a number of subgroups, which is where the "Pruner" class comes into play.
  • "Pruner": To receive a predefined number of subgroups for arbitrary RP algorithms, we follow a two-stage process: Grow a large tree by tweaking tree-specific hyperparameters and then prune it back to receive the desired number of subgroups. A "Partitioning" subclass is implemented such that it can first grow a large tree, for instance, with a low complexity parameter for rpart. Then, "Pruner" iteratively computes the SD of FMEs for each parent node of the current terminal nodes and removes all terminal nodes of the parent with the lowest SD.
  • "PartitioningPlot": Decouples visualizations of separating $\mathcal{D}$ into subgroups from the concrete implementation of the "Partitioning" subclass. Here, we make use of a dependency on partykit for a tree data structure. This allows visualizations of any partitioning with the same methods. The package ggparty (Borkovec and Madin, 2019) creates tree figures that illustrate the partitioning, descriptive statistics for each terminal node, and histograms of FMEs (and optionally NLM values).

### v2

@@ Introduction
Let us take a moment to compare the FME plot with the combined ICE and PD plot generated by the R package iml (Molnar et al., 2018) (see Fig. 4). This is one of the most popular and established model-agnostic ways to interpret predictive models (Molnar, 2022). The ICE is a local model explanation and represents the prediction for an observation where only the features of interest are varied (in this case, only ‘temp’). The PD is the average of ICEs (in the univariate case, the vertical average) and indicates the global, average prediction when a subset of features is varied for all observations. Although we can see a rough trajectory of the feature influence on local and average predictions, it is difficult to pinpoint the exact effects of changing ‘temp’ on the prediction for single observations. Furthermore, ICE curves are more likely to be subject to model extrapolations, a result of predicting in areas where the model was not trained on a sufficient amount of data.
@@
Figure 4: An ICE and PD plot for feature ‘temp’ generated by the R package iml. Each solid blue curve (an ICE) represents predictions for a single instance while only ‘temp’ varies. The dashed black curve (the PD) is the vertical average of ICEs and represents the average, isolated influence of ‘temp’.
FMEs allow for positive or negative step sizes. For instance, let us investigate the effects of an isolated drop in humidity by 10 percentage points. We can observe an AME of 108 additional predicted bike rentals a day. Individual effects tend to be larger the higher the humidity on that particular day.
@@
> effects.univariate.humidity = fme(+   model = forest,+   data = bikes,+   features = list("humidity" = -0.1),+   ep.method = "envelope")> summary(effects.univariate.humidity)
@@
Forward Marginal Effects ObjectStep type:  numericalFeatures & step lengths:  humidity, -0.1Extrapolation point detection:  envelope, EPs: 1 of 731 obs. (0 %)Average Marginal Effect (AME):  108.0477
> plot(effects.univariate.humidity)
Figure 5: Univariate FMEs for a drop in humidity by 10 percentage points. Especially for high humidity values, the drop results in a considerable increase in predicted daily bike rentals.
In many applications, we are interested in interactions of features on the prediction. Until now, we only analyzed the univariate effects of ‘temp’ and ‘humidity’ on the predicted amount of bike rentals. However, potential interactions between features may exist. We evaluate an increase in temperature by 5 degrees Celsius and a simultaneous drop in humidity by 10 percentage points (see Fig. 6). For a bivariate change in feature values, the two arrows depict the direction and magnitude of the feature change in the respective variable. As in the univariate case, we plot local averages within hexagons to avoid overplotting values. The location of the hexagon is determined by the observations’ observed feature values in the provided data set. Its color indicates the FME associated with the bivariate feature change. An increase in the outside temperature by 5 degrees Celsius and a simultaneous drop in humidity by 10 percentage points is associated with an AME of 414. The univariate AMEs roughly add up to the bivariate AME, indicating that, on average, there is no additional interaction between both feature changes on the prediction.
> effects.bivariate = fme(+   model = forest,+   data = bikes,+   features = list("temp" = 5, "humidity" = -0.1),+   ep.method = "envelope")> summary(effects.bivariate)
Forward Marginal Effects ObjectStep type:  numericalFeatures & step lengths:  temp, 5  humidity, -0.1Extrapolation point detection:  envelope, EPs: 49 of 731 obs. (7 %)Average Marginal Effect (AME):  413.6163
@@
Figure 6: Visualizing bivariate FMEs for an increase in ‘temp’ by 5 degrees Celsius and a simultaneous drop in ‘humidity’ by 10 percentage points. FMEs are highly heterogeneous. We can see mostly positive effects, especially for observations with combinations of medium ‘temp’ and ‘humidity’ values.
@@
Let us repeat the same procedure as for univariate feature changes and compare the FME plot to an alternative option, the bivariate PD plot (see Fig. 7). As opposed to the novel visualization with FMEs, the PD plot only visualizes the average, global effect of changing both features on the predicted amount of bike rentals. It does not inform us about the distribution of observed feature values, thus not allowing us to evaluate the effects of increasing one feature and decreasing another simultaneously.
@@
Figure 7: A bivariate PD plot (created via the R package iml), visualizing the global interaction between ‘temp’ and ‘humidity’ on the predicted amount of bike rentals. Plugging in medium to large values for ‘temp’ and low to medium values for ‘humidity’, ceteris paribus, results in more predicted bike rentals on average. As opposed to bivariate FMEs, we cannot investigate multiple local effects, nor can we see the actual distribution of observed feature values. As a result, we cannot evaluate the effects of increasing one feature and decreasing another simultaneously.
@@
Let us now proceed to investigate non-linearity. Non-linearity can be visually assessed for ICE curves (see Fig. 4), but it is hard to quantify and would be somewhat meaningless for a large variation in the feature of interest. Furthermore, for bivariate or higher-dimensional changes in feature values, we lose any option for visual diagnoses of non-linearity. In contrast, the NLM can be computed for FMEs with continuous step sizes, regardless of dimensionality. The average non-linearity measure (ANLM) is 0.36, indicating that the linear secant, on average, is a bad descriptor of the FME.
@@
> effects.bivariate.nlm = fme(+   model = forest,+   data = bikes,+   features = list("temp" = 5, "humidity" = -0.1),+   ep.method = "envelope",+   compute.nlm = TRUE)> effects.bivariate.nlmForward Marginal Effects ObjectFeatures & step lengths:  temp, 5  humidity, -0.1Average Marginal Effect (AME):  413.6163Average Non-Linearity Measure (ANLM):  0.36> plot(effects.bivariate.nlm, with.nlm = TRUE)
Figure 8: Adding NLM computations to the FME plot. Each hexagon in the left and right plots represents a local average of FME and NLM values, respectively.
Fig. 8 simply contrasts FME values with the corresponding NLM values. In this case, we can see both non-linear FMEs (whiter NLM) and linear FMEs (deep blue-colored NLM). We could now, for instance, focus on interpreting linear FMEs. All FMEs depicted in Fig. 9 have an NLM of 0.9 or higher, meaning that they almost fully describe the model prediction for proportional changes in ‘temp’ and ‘humidity’.
Figure 9: Visualizing FMEs with an NLM $\geq$ 0.9.
An advantage of FMEs is their ability to provide comprehensible model insight even when exploring higher-order feature changes. Let us factor in a third feature change, now simultaneously reducing windspeed by 5 miles per hour, and visualize the distribution of FME and NLM values. We can see that in addition to an increase in temperature and a decrease in humidity, a decrease in windspeed further boosts the average number of predicted daily bike rentals.
> effects.trivariate.nlm = fme(+   model = forest,+   data = bikes,+   features = list("temp" = 5, "humidity" = -0.1, "windspeed" = -5),+   ep.method = "envelope",+   compute.nlm = TRUE)> summary(effects.trivariate.nlm)Forward Marginal Effects ObjectStep type:  numericalFeatures & step lengths:  temp, 5  humidity, -0.1  windspeed, -5Extrapolation point detection:  envelope, EPs: 117 of 731 obs. (16 %)Average Marginal Effect (AME):  537.7385Average Non-Linearity Measure (ANLM):  0.33> plot(effects.trivariate.nlm, with.nlm = TRUE)
Figure 10: Adding a third feature change, a drop in windspeed by 5 miles per hour, and visualizing the distribution of FME and NLM values. For the NLM plot, negative NLMs are binned as 0. It follows that the ANLM value in the plot differs from the raw ANLM in the summary output.
@@
So far, we have only evaluated changes in continuous features. In many applications, we are concerned with switching categories of categorical features, a way of counterfactual thinking inherent to the human thought process. Note that despite the allure of switching categories of categorical features, one needs to be aware of potential model extrapolations. To illustrate this, we switch each non-rainy day’s precipitation status to rainfall. Rainfall has an average isolated effect of lowering daily rentals by 803 bikes (see Fig. 11).
@@
> effects.categ = fme(+   model = forest,+   data = bikes,+   features = list("weather" = "rain"))> summary(effects.categ)Forward Marginal Effects ObjectStep type:  categoricalFeature & reference category:  weather, rainExtrapolation point detection:  none, EPs: 0 of 710 obs. (0 %)Average Marginal Effect (AME):  -802.8716> plot(effects.categ)
@@
Figure 11: Distribution of categorical FMEs resulting from switching each non-rainy day’s precipitation status to rain. On average, rainfall lowers predicted bike rentals by 803 bikes per day.
@@
In our examples, we can see highly heterogeneous local effects. The more heterogeneous FMEs are, the less information the AME carries. In many practical applications, we are interested in compactly describing the behavior of the predictive model across the feature space, akin to a beta coefficient in a linear model. This is where regional explanations come into play. We aim to find subgroups with more homogeneous FME values, thereby describing the behavior of the model not in terms of a global average but in terms of multiple regional averages (cAMEs).
In fmeffects, this can be achieved by further processing the ForwardMarginalEffect object containing FMEs (and optionally NLM values) using the came() function. This returns a Partitioning object (in this case, an object of the class "PartitioningCTREE", a subclass of the abstract class "Partitioning", see later section on Design and options for extensions).
For the univariate change in temperature by 5 degrees Celsius, we decide to search for precisely 2 subgroups [^2] This value is to be set by the user depending on how many regional explanations are to be found. Alternatively, we can search for a pre-defined SD of FMEs inside the terminal nodes. How many subgroups can be found depends on the data and predictive model. (for a description of this algorithm, see the following section on Design and options for extensions). A summary of the created object informs us about the number of observations, cAME, and standard deviation (SD) of FMEs inside the root node and leaf nodes (the found subgroups). We succeeded in finding subgroups with lower SDs while maintaining an appropriate sample size. The root node SD of 611 can be successfully split down to 437 and 355 within the subgroups. By visualizing the tree, we can see how the data was partitioned. For cooler outside temperatures equal to or lower than $\approx$ 16 degrees Celsius, we can observe a positive cAME of 728 additional bike rentals per day. On warmer days with a temperature above $\approx$ 16 degrees Celsius, the model predicts 196 less bike rentals a day when the outside temperature increases by 5 degrees.
> subspaces = came(effects = effects.univariate.temp, number.partitions = 2)> summary(subspaces)
PartitioningCtree of an FME objectMethod:  partitions = 2   n      cAME  SD(fME) 683  307.3275 611.0778 * 372  728.3942 437.0463 311 -196.3278 354.5090---* root node (non-partitioned)AME (Global): 307.3275
@@
Figure 12: Using a decision tree to find subgroups of observations with more homogeneous FMEs of increasing ‘temp’ by 5 degrees Celsius. Each leaf node visualizes one subgroup, the number of observations, the cAME, and the SD of FMEs indicating FME homogeneity.
@@
When to search for regional explanations thus depends on the heterogeneity of local effects. The ame() function provides an appropriate summary for the entire model. It uses a default step size of 1 or 0.01 for small feature ranges. For categorical FMEs, it uses every observed category as a reference category. Alternatively, custom step sizes and subsets of features can be used. The summary() function prints a compact model summary of each feature, a default step size, the AME, the SD of FMEs, 25% and 75% quantiles of FMEs, as well as the number of observations left after excluding extrapolation points (EPs). A large dispersion indicates heterogeneity of FMEs and thus a small fidelity of the AME and possible benefits from searching for subgroups with varying cAMEs. A different workflow can, therefore, also consist of starting with the table generated by ame() and deciding which feature effects can be described by AMEs and which might be better describable by subgroups and cAMEs. If this has been unsuccessful, we can resort to local model explanations. Recall our example from the previous section on Regional explanations where we split FMEs associated with increasing temperature by 5 degrees Celsius. From the ame() summary, we see that ‘temp’ has a relatively large SD in relation to its AME (here calculated with a step size of 1), and the interquartile range indicates a wide spread of FMEs from -20 in the 25% quantile up to 108 in the 75% quantile, which makes it a promising candidate to find subgroups with more homogeneous FMEs.
> ame.results = ame(model = forest, data = bikes)> summary(ame.results)
@@
Model Summary Using Average Marginal Effects:      Feature step.size        AME       SD       0.25       0.75   n1      season    winter  -942.0906 466.3691 -1298.1011  -617.5663 5502      season    spring   136.2185 569.5307  -244.4237   650.0125 5473      season    summer   293.6264 549.2972   -42.7551   738.2056 5434      season      fall   533.5502 579.5541    52.3706  1138.0863 5535        year         0 -1899.4966 639.1695 -2354.1389 -1506.0582 3666        year         1  1790.6269 524.4711  1421.7925  2194.1396 3657     holiday        no     195.93  218.386   123.2468   228.0909  218     holiday       yes  -133.3134 154.8869  -201.3635   -25.1245 7109     weekday    Sunday   155.5219 188.8708     9.3486   252.0308 62610    weekday    Monday  -158.9218 215.5047  -263.2441    -4.8485 62611    weekday   Tuesday  -115.7316 193.4508  -197.7396    13.3208 62612    weekday Wednesday   -44.3056 173.8664  -115.5562    63.1344 62713    weekday  Thursday     16.005  161.125   -61.1673    89.5043 62714    weekday    Friday    57.1498 163.5602   -27.6519    128.752 62715    weekday  Saturday   103.7648 170.5678    -0.2044    178.493 62716 workingday        no   -42.8794 139.8572  -145.7104    66.2131 50017 workingday       yes    48.1298 158.3666   -60.2448   145.5003 23118    weather     misty  -221.5664 328.3458  -413.4363   -69.4238 48419    weather     clear   385.8674 347.6119   162.2048   476.8631 26820    weather      rain  -802.8716 384.2624 -1054.7158  -543.2614 71021       temp         1    58.0487 164.8714   -20.0019   108.4669 73122   humidity      0.01     -19.86  62.1753   -36.5407    10.4535 73123  windspeed         1   -24.7315  77.1757   -56.9247    13.7468 731
@@
The fmeffects package is built on a modular design for improved maintainability and future extensions. Fig. 13 provides a visual overview of the core design. The greatest emphasis is placed on the strategy and adapter design patterns (Gamma et al., 1994). Simply put, the strategy pattern decouples the source code for algorithm selection at runtime into separate classes. We repeatedly implement this pattern throughout the package by creating abstract classes whose subclasses implement specific functionalities. The adapter design pattern (also called a “wrapper”) creates an interface for communication between two classes.
@@
  • "Predictor": An abstract class that implements the adapter pattern to accommodate future implementations of storing a predictive model. "PredictorMLR3", "PredictorParsnip", and "PredictorCaret" are subclasses that store an mlr3, parsnip (Kuhn and Vaughan, 2023) (part of tidymodels), or caret model object. This allows users of fmeffects to use numerous predictive models such as random forests, gradient boosting, support vector machines, or neural networks. "PredictorLM" stores models returned by lm(), glm(), or gam(). The package can be extended with novel model types by implementing a new subclass that stores the model, data, target, and is able to return predictions.
  • "AverageMarginalEffects": A class to compute AMEs for each feature in the data (or a subset of features). Internally, a new "ForwardMarginalEffect" object is used to compute and aggregate FMEs. For convenience, we implement a wrapper function ame() to facilitate object creation and to initiate computations without requiring user input in the form of R6 syntax.
  • "ForwardMarginalEffect": The centerpiece class of the package. It keeps access to a Predictor, stores important information to create FMEs, and after the computations are completed, stores results and gives access to visualization methods. For convenience, the wrapper function fme() can be used.
  • "FMEPlot": An abstract class for code decoupling of different plot categories into distinct classes. Subclasses include "FMEPlotUnivariate", "FMEPlotBivariate", "FMEPlotHigherOrder",
    "FMEPlotCategorical".
  • "ExtrapolationDetector": Identifies (and excludes) EPs. The current implementation supports the method “envelope”, excluding points outside the multivariate envelope of the training data.
  • "NonLinearityMeasure": For the NLM, we need to approximate three line integrals, e.g., via Simpson’s 3/8 rule. The general definition of Simpson’s 3/8 rule for a univariate function f(x) and integration interval $[a,b]$ corresponds to:
@@
  • "Partitioning": An abstract class, allowing for various implementations of finding subgroups for cAMEs. For convenience, the wrapper function came() can be used. The current implementation supports RP via the rpart and partykit (CTREE algorithm) packages (classes "PartitioningRPart" and "PartitioningCTREE").
    We believe there are two criteria that should guide this process: FME homogeneity within each subgroup and the number of subgroups. A low number of subgroups is generally preferred. In certain applications, we may want to search for a predefined number of subgroups, akin to the search for a predefined number of clusters in clustering problems. Many RP algorithms do not support searching for a number of subgroups, which is what the "Pruner" class is intended for.
  • "Pruner": To receive a predefined number of subgroups for arbitrary RP algorithms, we follow a two-stage process: grow a large tree by tweaking tree-specific hyperparameters and then prune it back to receive the desired number of subgroups. A "Partitioning" subclass is implemented such that it can first grow a large tree, e.g., with a low complexity parameter for rpart. Then "Pruner" iteratively computes the SD of FMEs for each parent node of the current terminal nodes and removes all terminal nodes of the parent with the lowest SD.
  • "PartitioningPlot": Decouples visualizations of the separation of $\mathcal{D}$ into subgroups from specific implementations of the "Partitioning" subclass. Here, we make use of a dependency on partykit for a tree data structure. This allows visualizations of any partitioning with the same methods. The package ggparty (Borkovec and Madin, 2019) creates tree figures that illustrate the partitioning, descriptive statistics for each terminal node, and histograms of FMEs (and optionally NLM values).